OpenStudy (anonymous):

Sigma r=1 to r=7 tan square {r(pi)/16} Equals to what?

Parth (parthkohli):

Hmm, I'm getting 35...

OpenStudy (anonymous):

Yup how?

Parth (parthkohli):

Have to do a lot of typing for that...

Parth (parthkohli):

OK, I'll type.\[\tan^2\frac{\pi}{16} + \tan^2 \frac{2\pi}{16} + \tan^2\frac{3\pi}{16} + \tan^2\frac{4\pi}{16} + \cot^2 \frac{3\pi}{16} + \cot^2 \frac{2\pi}{16} + \cot^2 \frac{\pi}{16}\]

Parth (parthkohli):

Now \(\tan^2 4\pi/16 = 1\). We'll pair the rest like this:\[\tan^2 \pi/16 + \cot^2 \pi/16= \left(\tan \pi/16 + \cot \pi/16\right)^2 - 2\]\[= \left(\frac{1}{\sin \pi/16 \cos \pi/16}\right)^2 -2\]

Parth (parthkohli):

In the end we have\[4\left(\frac{1}{\sin^2 \pi/8} + \frac{1}{\sin^2 \pi/4 } + \frac{1}{\sin ^2 3\pi/8 }\right)-5\]Now again notice \(\sin^2 3\pi/8 = \cos^2 \pi/8\)

Parth (parthkohli):

Sorry, this website is very glitchy at the moment. Can you do the rest of the work on your own? Thanks :)

OpenStudy (anonymous):

Yaa..sure thanks a lot..got the method.

Parth (parthkohli):

Great, are you in 11th?

OpenStudy (anonymous):

12th